Solve the following polynomial equation. If needed, write your answer as a fraction reduced to lowest terms.
y^4 + 8y^2 -9 = 0

I'd like to know the steps to completing this. Can this be factored? How do I make it ready to be plugged into the quadratic formula? Any help is appreciated. Thanks in advance.

Answer by stanbon(75887)

y^4 + 8y^2 -9 = 0
-----
Factor:
y^4 + 9y^2-y^2 -9 = 0
----------------------------
y^2(y^2+9) -(y^2+9) = 0
----------------------------
Factor:
(y^2+9)(y^2-1) = 0
---
Factor:
(y^2+9)(y-1)(y+1) = 0
---
y = 3i or y = -3i or y = 1 or y = -1
=======================================
Cheers,
Stan H.
=============